# Automatic Input Feature Relevance via Spectral Neural Networks

### v1

@@ Abstract
Working with high-dimensional data is a common practice, in the field of machine learning. Identifying relevant input features is thus crucial, so as to obtain compact dataset more prone for effective numerical handling. Further, by isolating pivotal elements that form the basis of decision making, one can contribute to elaborate on - ex post - models’ interpretability, so far rather elusive. Here, we propose a novel method to estimate the relative importance of the input components for a Deep Neural Network. This is achieved by leveraging on a spectral re-parametrization of the optimization process. Eigenvalues associated to input nodes provide in fact a robust proxy to gauge the relevance of the supplied entry features. Unlike existing techniques, the spectral features ranking is carried out automatically, as a byproduct of the network training. The technique is successfully challenged against both synthetic and real data.
preprint: APS/123-QED
@@
Deep Neural networks (DNNs) [1, 2, 3, 4, 5] are among the most popular machine learning tools, and, as such, routinely employed in a vast plethora of applications [6, 7, 8, 9]. As for other machine learning techniques, and when employed for e.g. classification and/or regression tasks, DNNs can automatically extract valuable information from the supplied input vector. This amounts to isolate crucial components that prove decisive in implementing unmanned decision making strategies. In many cases, however, input data are naturally embedded in high dimensional spaces. This curse of dimensionality, quoting with lexical abusing from a renowned motto [10], makes it in general difficult to selectively single out, from the sea of underlying components, the subset of attributes deemed essential. Despite this degree of inherent complexity, it is of the utmost relevance to identify key input features that could be instrumental in (i) devising apt solutions for data compression and compactification; (ii) providing critical insights into the processes that led to the produced output response.
A first solution is to operate an a priori features’ selection, with the aim to isolate beforehand a limited subset of supposedly relevant input elements that should undergo subsequent scrutiny [11]. This certainly yields machine learning models with a reduced dimensional load. Prior trimming criteria could however resolve in the removal of essential information, erroneously targeted as fruitless by the upstream selection.
@@
Other methods have been proposed for post-hoc interpretability, that is, to shed light on the pathways that form the basis of decision of a trained model [12, 13, 14, 15, 14, 16, 17]. Worth mention is the so called Layer-Wise Relevance Propagation method [16, 15] which implements a backprogation, from the output to the input space, to reach the relevant features that populate the entry layer. The method has been successfully challenged against distinct application realms [18, 19]. However, like many other approaches belonging to the same class, the latter is an input-dependent framework and requires a certain number of operations to be carried out after training, for each examined input, so as to assess the relevance of the features for that particular supplied item.
Here, we put forward a different method with complementary objectives. The aim is to define a strategy to quantify the relative importance of different input feature, for a dataset taken as a whole, and for a general feedforward neural network. Unlike the aforementioned methods, the proposed metric is not input dependent. At variance, it bears universality traits, by providing a global insight on how the model oversees the input space, upon learning. Moreover, unlike standard techniques for input feature selection, the analysis is not conducted a priori on the data to be eventually scanned. The ranking of the features is indeed carried out automatically, as a byproduct of the training stage. This is achieved by establishing an ideal bridge, justified on first principles, between a pool of trainable parameters – to be optimized for task resolution – and the level of inherent relevance, as associated to different components of the input vector. At the end of the training, by accessing the values assigned to the aimed parameters, will therefore yield an automatic and self-consistent ranking of the individual input features.
@@
To this end, we will make use of a recently introduced scheme for handling the optimization of feedforward networks, known as spectral parametrization [20, 21, 22, 23, 24, 25, 26]. The mathematical foundation of the spectral methodology lies within the theory of networks. From this angle, feedforward neural networks can be pictured via a dedicated spectral decomposition that materializes, in direct space, in a collection of adjacency matrices for the linear mapping of the signal across adjacent layers.
@@
Let us begin by reviewing the basic of the spectral method [20]. We denote with $l-1$ and $l$ two consecutive layers of a feedforward neural network. Assume the layers to be respectively composed by $N$ and $M$ nodes. The vector $\textbf{x}^{l-1}=\left(x^{l-1}_{1},\ldots,x^{l-1}_{N}\right)$ (resp. $\textbf{x}^{l}=\left(x^{l}_{1},\ldots,x^{l}_{M}\right)$) stores information on the activity of the nodes belonging to layer $l-1$ (resp. $l$). The linear transfer between the two adjacent layers is operated by a rectangular $M\times N$ matrix $W$, whose elements read $w_{ji}$. Hence, $\textbf{x}^{l}=f(W\textbf{x}^{l-1})$, where $f$ is a non-linear filter. Under the spectral paradigm, the above operation can be recast in an equivalent form that assumes dealing with the bipartite direct network made of the $N+M$ nodes involved in the transfer. The state vector at the departure layer can be written as $\textbf{z}^{l-1}=\textbf{x}^{l-1}=\left(x^{l-1}_{1},x^{l-1}_{2},\ldots x^{l-1}_{N},0,0,\ldots 0\right)$ where we made explicit that the non trivial activity is solely localized on the nodes of the first layer. Then, $\textbf{z}^{l}=f(A\textbf{z}^{l-1})$, where the last $M$ elements of vector $\textbf{z}^{l}$ refer to the activity on the destination layer; $A$ is a $(M+N)\times(M+N)$ matrix with a $M\times N$ sub-diagonal block that coincides matrix $W$, as depicted in the annexed Appendix A. It should be noted that possible diagonal entries of matrix $A$ do not reflect on the produced activity on layer $l$. One can invoke therefore a spectral decomposition of the square transfer matrix, as $A=\Phi\Lambda\Phi^{-1}$ where (i) $\Phi$, the matrix of the eigenvectors, is lower-block triangular with ones on the diagonal, (ii) $\Lambda$ identifies the diagonal matrix formed with the eigenvalues ($\lambda_{i}$, $i=1,\ldots N+M$) of $A$. Given the specific form of matrix $\Phi$, it can be shown [24, 21] that $\Phi^{-1}=2\mathbb{I}-\Phi$. In practical terms, the elements of $A$, namely those involved in the self-consistent definition of the relevant transfer matrix $W$, can be written as a function of the eigenvalues $\lambda_{i}$ and the non trivial entries $\psi_{ji}$ of the eigenvector matrix $\Phi$. An explicit calculation can be performed that yields $w_{i\rightarrow j}=w_{ji}=(\lambda_{i}-\lambda_{j})\psi_{ji}$, where $i=1,...,N$ refer to nodes located on layer $l-1$ and $j=1,...,M$ to those assigned to layer $l$. Further technical details are provided in Appendix A. A direct correspondence can be hence drawn between individual eigenvalues and nodes within the sampled collection. Training a feedforward network under the spectral standpoint amounts to minimize the loss (a suitable metric that gauges the distance between produced and expected output at the exit layer) with respect to the eigenvalues and eigenvectors. In light of the existing correspondence, the magnitude of the optimized eigenvalues can be used as a reliable marker of the nodes’ relevance  [^1] The usage of the eigenvalues as an indirect measure of the nodes relevance has been explored in [24, 21, 25]. An effective trimming strategy was in particular proposed in [24] which enables for the trained network to be significantly reduced, ex post, by cutting nodes flagged as unimportant. By leveraging on the eigenvalues and their associated magnitudes, it was possible in [25] to estimate the hidden dimension of the target objective function under analysis., an observation that can be exploited for proposing and effective (input) feature detection algorithm, as we shall set to demonstrate in the following. To this end, we chose to operate under the simplifying assumption $\lambda_{j}=0$ (i.e. by forcing to zero the eigenvalues associated to destination nodes) yielding:
@@
The spectral writing of inter-nodes weights $w_{ji}$ introduced a multiplicative factor, the eigenvalue $\lambda_{i}$, shared by all links emanating from the departure node $i$. The larger the magnitude of the optimized $\lambda_{i}$ the more relevant the nodes for the processing of the information across the network. Specifically, we will convincingly conclude on the ability of the eigenvalues associated to the starting layer $l=1$ to automatically discriminate between irrelevant and relevant input features. Further, we will show that eigenvalues (multiplied by the norm of the corresponding eigenvector entry) can be exploited to rank the relative importance of the features otherwise deemed as irrelevant. The method will be tested against classification problems constructed from mock and real datasets. For all examined cases, a penalty is added to the loss function as in the spirit of [25]. This is a $L_{2}$ regularization of $\lambda_{i}$ which forces eigenvalues associated with irrelevant components to be set to zero (and not just left unaltered to their initial values, possibly drowned in the sea of their relevant homologues). It is worth stressing that in principle it is enough to have just the first layer parameterized as spectral for this analysis to apply. For further details on the training procedure refers to Appendix B.
@@
Independent Gaussian distributions dataset The first dataset is structured with 2 target classes and input vectors with 20 independent features. Each data component is sampled from a Gaussian distribution. While some components are drawn from distributions which are set to be identical across classes, others display only partial overlap. Components with identical distributions offer no discriminatory information. As such, they prove ineffective for classification purposes. Conversely, non trivial components are pivotal, the smaller the overlap the larger their relevance in promoting the sought discrimination.
@@
A spectral feedforward neural network, parametrized as stipulated by Eq. (1), has been trained with the aim of assigning input from the above dataset to the correct class of origin. The employed network consists of three hidden layers with, respectively, 100, 100, and 50 neurons. In addition, the network accommodates for an input layer of size 20 and an output layer of size 2. At the end of the training, we analyzed the 20 eigenvalues, associated with the input layer as in the spirit of the above discussion (with an abuse of language we will from here call eigenvalues the rescaled version of this latter which incorporate the norm of the eigenvectors, as stipulated in Appendix A). As shown in panel (a) of Fig. 1, eigenvalues entangled with irrelevant features are all zero, while eigenvalues which can be traced back to relevant input components are non-zero. Hence, the eigenvalues interfused with the input layer can efficiently tag features identified as relevant by the network itself to solve the assigned task. The magnitude of non-zero eigenvalues, i.e., those associated with relevant components, is found to correlate with a direct measure of the relative distance between the distributions from which individual components are drawn (see panel (b) of Fig. 1). Since the distributions differ only in their mean value, we quantify the above distance as the absolute value of the difference between the two means.
@@
Figure 1: Panel (a) displays the 20 post-training eigenvalues associated to the input layer, for the simple Gaussian dataset. The red dots and the red bars correspond to the average values and the standard deviations of the results obtained by repeating the experiments for 10 independent realizations. In panel (b) the same post-training eigenvalues are reported against a relevance parameter that gauges the distance between the distributions (respectively blue and red, in Fig. 5) from which the components are eventually drawn. The distance is defined as the absolute value of the difference between the two means.
@@
Correlated Gaussian distributions dataset Unlike the dataset just analyzed, real datasets often exhibit correlations between input vector components, and crucial task-solving information are frequently concealed within those correlations. With this motivation in mind, we tested the proposed spectral feature detection method on a different dataset that contains pairwise correlations between input vector components. The dataset is still organized into two distinct classes that are to be eventually identified. Input vectors are now made of 10 components. The first 8 components are defined as follows:
$\begin{split}&x_{2n}\sim\mathcal{N}(0,1)\\ &x_{2n+1}=x_{2n}\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\text{if}\;\;c(\vec{x})=0\\ &x_{2n+1}=\begin{cases}x_{2n}\;\text{with prob.}\;1-p\\ -x_{2n}\;\text{with prob.}\;p\end{cases}\;\;\text{if}\;\;c(\vec{x})=1\end{split}$ (2)
@@
where $c(\vec{x})$ is the target class of the input and $n\in\{0,1,2,3\}$. The last two components are independently drawn from two identical normal distributions $N(1,0)$ regardless of the class of belonging:
@@
As a first noticeable observation, we remark that it is not possible to separate the two datasets by solely relying on independent observations of the respective vector components. Indeed, crucial information are reflected in how pairs of components are mutually correlated.
@@
To better understand the relationship between the different components, let us focus on the first two entries of the input vectors, $x_{0}$ and $x_{1}$. For data belonging to class 00 (as we label the first of the two) $x_{0}=x_{1}$, that is, the first two components are perfectly correlated. Conversely, the two components of data belonging to the other class (denoted by $1$) are equal with probability $1-p$ and have opposite sign with probability $p$. If $p=0$, the first two components are equal for both classes. No information about the class to which the supplied data belong can be thus extracted by comparing the two above components. The second component is indeed a mere copy of the first and therefore irrelevant for classification purposes. Conversely, if $p=1$, the first and second components from class $1$ data are in perfect anti-correlation, while data from class 00 are still matched and thus fully correlated. Hence, in the latter case, it is possible to completely separate the two classes by solely comparing $x_{0}$ and $x_{1}$. Parameter $p$ quantifies therefore the degree of sensible information which can be pulled out from a direct comparison of the two aforementioned components. Finally, the last two components are not correlated to the others and identically distributed for both classes, hence devoid of information. We hence assign a parameter $p=0$ to these two components to explicitly state their manifest irrelevance.
@@
Figure 2: Panel (a) reports the ten post-training eigenvalues (refereed to the input layer) for the correlated Gaussian dataset.Red dots and red bars quantifies the recorded average values and the associated standard deviations, as obtained by averaging over 10 independent realizations. In panel (b) the same post-training eigenvalues are plotted against the parameter $p$, as defined in the main text
@@
The architecture of the employed neural network, parametrized via its spectral modality, is identical to that put forward for the analysis of the uncorrelated dataset (except for the first layer that in this case receives just 10 entries). In panel (a) of Fig. 2, the post-training eigenvalues referred to the input layer are shown. Firstly, we notice that pairs are correctly spotted out, as they are assigned the very same eigenvalue entry. Moreover, eigenvalues associated with irrelevant components are tuned zero for this dataset as well, an observation that testifies again on the ability of the spectral network to discriminate between irrelevant and relevant features. Lastly, the magnitude of the recorded eigenvalues scales proportionally to the control parameter $p$, which quantifies the relevance of the selected pair for classification purposes (panel (b), Fig. 2). In conclusion, and also for this second, more complex dataset, the optimized eigenvalues referred to the first layer are a good proxy of the importance of the associated components.

### v2

@@ Abstract
In machine learning practice it is often useful to identify relevant input features, so as to obtain compact dataset for more efficient numerical handling. On the other hand, by isolating key input elements, ranked according their respective degree of relevance, can help to elaborate on the process of decision making. Here, we propose a novel method to estimate the relative importance of the input components for a Deep Neural Network. This is achieved by leveraging on a spectral re-parametrization of the optimization process. Eigenvalues associated to input nodes provide in fact a robust proxy to gauge the relevance of the supplied entry features. Notably, the spectral features ranking is performed automatically, as a byproduct of the network training, with no additional processing to be carried out. The technique is successfully challenged against both synthetic and real data.
1 Introduction
@@
Deep Neural networks (DNNs) [1, 2, 3, 4, 5] are among the most popular machine learning tools, and, as such, routinely employed in a vast plethora of applications [6, 7, 8, 9]. As for other machine learning techniques, and when employed for e.g. classification and/or regression tasks, DNNs can automatically extract valuable information from the supplied input vector. This amounts to isolate crucial components to implement unmanned decision making strategies. In many cases of interest, input data are multidimensional. In general, it is hence difficult to selectively identify the subset of attributes which prove crucial for the analysis. To identify key input features could be indeed relevant for (i) devising apt solutions for data compression and compactification; (ii) providing critical insights into the processes that led to the produced output.
A first solution is to operate an a priori feature selection, with the aim to isolate beforehand a limited subset of supposedly relevant input elements that should undergo subsequent scrutiny [10]. This is certainly useful to reduce the size of the input to be eventually processed. On the other hand, by clipping part of the dataset, a non zero risk always exists that could yield an unpredictable loss of performance for the trained network. This is as opposed to other viable approaches designed so as to preserve the total amount of information across training.
@@
Among the methods proposed to shed light on the pathways that form the basis of decision of a trained model, worth mentioning is the so called Layer-Wise Relevance Propagation method [11, 12]. This latter implements a backprogation, from the output to the input space, to reach the relevant features that are provided at the entry layer. The method has been successfully challenged against distinct domains of application [13, 14]. Reference should also be made to SHAP [15] (SHapley Additive exPlanations) a game theoretic approach based on the classic Shapley values, which connects optimal credit allocation to local output explanation. Other examples include - but are not limited to - the techniques identified via the compact acronyms IG, DeepLIFT, LIME [16, 17, 18, 19, 20, 21]. To the best of our knowledge, the vast majority of existing methods are however input-dependent. As such, they require a certain number of operations to be carried out after training, for each examined input, so as to assess the relevance of the features for any given supplied item.
Here, we put forward a different method which yields a global metric to assess the relevance of input features to a general feedforward neural network. The procedure applies to the whole dataset and not just to individually examined items. Furthermore, the evaluation of the features is obtained as a byproduct of the training with no additional processing to be carried out. At the end of the training a ranking of the supplied feature is returned, as determined by a set of scalar parameters (one for each distinct feature provided as an input). As mentioned, and unlike the aforementioned methods, the proposed metric is not input dependent. At variance, it provides a global insight on how the model oversees the input space, upon learning. Also, and as opposed to standard techniques for input feature selection, the analysis is not conducted a priori on the data to be eventually scanned.
@@
To reach this goal, we will make use of a recently introduced scheme for handling the optimization of feedforward networks, known as spectral parametrization [22, 23, 24, 25, 26, 27]. The mathematical foundation of the spectral methodology lies within the theory of networks. From this angle, feedforward neural networks can be pictured via a dedicated spectral decomposition that materializes, in direct space, in a collection of adjacency matrices for the linear mapping of the signal across adjacent layers. As we shall argue, the optimized eigenvalues, as made accessible at the end of the training, will return an automatic and self-consistent ranking of the individual input features.
The paper is organized as follows. In the next Section we will review the mathematical basis of the spectral approach to neural network training and introduce the propose ranking strategy. Then, we will turn to discussing a first application to independent and correlated Gaussian data. These are mock data, with different level of complexity, to be regressed by the spectral neural network. In the subsequent Section the proposed method is challenged against the celebrated MNIST database, operated in different modalities. As we will show, the spectral based feature detection method is able to isolate a subset of relevant pixels that prove central for image classification. Then, we turn to challenging the method against an additional dataset, made of stellar spectra. The goal here is to obtain information on relevant frequency bands for the assessment of stellar metallicity. Finally, we sum up and draw our conclusions.
2 On the spectral parametrization
@@
Let us begin by reviewing the basic of the spectral method [22]. We denote with $l-1$ and $l$ two consecutive layers of a feedforward neural network. Assume the layers to be respectively composed by $N$ and $M$ nodes. The vector $\textbf{x}^{l-1}=\left(x^{l-1}_{1},\ldots,x^{l-1}_{N}\right)$ (resp. $\textbf{x}^{l}=\left(x^{l}_{1},\ldots,x^{l}_{M}\right)$) stores information on the activity of the nodes belonging to layer $l-1$ (resp. $l$). The linear transfer between the two adjacent layers is operated by a rectangular $M\times N$ matrix $W$, whose elements read $w_{ij}$. Hence, $\textbf{x}^{l}=f(W\textbf{x}^{l-1})$, where $f$ is a non-linear filter. Under the spectral paradigm, the above operation can be recast in an equivalent form that assumes dealing with the bipartite direct network made of the $N+M$ nodes involved in the transfer. The state vector at the departure layer can be written as $\textbf{z}^{l-1}=\textbf{x}^{l-1}=\left(x^{l-1}_{1},x^{l-1}_{2},\ldots x^{l-1}_{N},0,0,\ldots 0\right)$ where we made explicit that the non trivial activity is solely localized on the nodes of the first layer. Then, $\textbf{z}^{l}=f(A\textbf{z}^{l-1})$, where the last $M$ elements of vector $\textbf{z}^{l}$ refer to the activity on the destination layer; $A$ is a $(M+N)\times(M+N)$ matrix with a $M\times N$ sub-diagonal block that coincides matrix $W$, as illustrated in the Methods section. It should be noted that the diagonal entries of matrix $A$ do not reflect on the produced activity on layer $l$, see Figure 1. One can hence invoke a spectral decomposition of the square transfer matrix, as $A=\Phi\Lambda\Phi^{-1}$ where (i) $\Phi$, the matrix of the eigenvectors, is lower-block triangular with ones on the diagonal, (ii) $\Lambda$ identifies the diagonal matrix formed with the eigenvalues ($\lambda_{i}$, $i=1,\ldots N+M$) of $A$. Given the specific form of matrix $\Phi$, it can be shown [26, 23] that $\Phi^{-1}=2\mathbb{I}-\Phi$. In practical terms, the elements of $A$, namely those involved in the self-consistent definition of the relevant transfer matrix $W$, can be written as a function of the eigenvalues $\lambda_{i}$ and the non trivial entries $\psi_{ij}$ of the eigenvector matrix $\Phi$. An explicit calculation can be performed that yields $w_{i\rightarrow j}=w_{ji}=(\lambda_{i}-\lambda_{j})\psi_{ji}$, where $i=1,...,N$ refer to nodes located on layer $l-1$ and $j=1,...,M$ to those assigned to layer $l$. Further technical details are provided in the Methods section. A direct correspondence can be hence drawn between individual eigenvalues and nodes within the sampled collection. Training a feedforward network under the spectral standpoint amounts to minimize the loss (a suitable metric that gauges the distance between produced and expected output at the exit layer) with respect to the eigenvalues and eigenvectors. In light of the existing correspondence, the magnitude of the optimized eigenvalues can be used as a reliable marker of the nodes’ relevance, an idea that has been explored in [26, 23, 27]. An effective trimming strategy was in particular proposed in [26] which enables for the trained network to be considerably shrunk in size by removing nodes flagged as unimportant. By leveraging on the eigenvalues and their associated magnitudes, it was possible in [27] to estimate the hidden dimension of the target objective function. By expanding further on these ideas, we will hereafter propose an effective input feature detection algorithm. To this end, we chose to operate under the simplifying assumption $\lambda_{j}=0$ (i.e. by forcing to zero the eigenvalues associated to destination nodes) yielding:
@@
The spectral rewriting of the inter-nodes weights $w_{ij}$ yielded therefore a factor, the eigenvalue $\lambda_{i}$, which multiplies all the links emanating from the departure node $i$. The larger the magnitude of the optimized $\lambda_{i}$ the more relevant the associated node $i$ for the processing of the information across the network. Building on this observation, we will show that the trained eigenvalues associated to the starting layer $l=1$ can be effectively employed to automatically discriminate between irrelevant and relevant input features. More specifically, we will demonstrate, that eigenvalues (multiplied by the norm of the corresponding eigenvector entry) can be exploited to rank the relative importance of the features otherwise deemed as irrelevant. The method will be tested against classification problems constructed from mock and real datasets. For all examined cases, a penalty is added to the loss function as in the spirit of [27, 28]. This is a $L_{2}$ regularization of $\lambda_{i}$ which forces eigenvalues associated with irrelevant components to be set to zero (and not just left unaltered to their initial values, possibly drowned in the sea of their relevant homologue). It is worth stressing that in principle it is enough to have just the first layer parameterized as spectral for this analysis to apply. For further details on the training procedure refers to the Methods section.
$l-1$$l$$l-1$$l$$l-1$$l$$l-1$$l$
$=$$A$$\textbf{z}^{l-1}$$\textbf{z}^{l}$$=$$A$$\textbf{z}^{l-1}$$\textbf{z}^{l}$
Figure 1: The linear information flow for a modified network with the inclusion of diagonal entries (eigenvalues) in the triangular matrix $A$. The eigenvalues introduce self-loops (as depicted on the right), and this is at variance with the usual setting without loops (on the left). Notice that the bottom elements of $\textbf{z}^{l}$ are identical in both cases, yielding exactly the same activation vector upon linear transfer
3 Numerical experiments
The method discussed above for input feature detection is hereafter challenged against different datasets. We will begin in particular by considering a simple multidimensional Gaussian dataset with independent entries. Then, we will turn to considering mock data generated from correlated Gaussian variables. After that, we will move forward to work with the celebrated MNIST dataset and, finally, we will focus on an experimental catalogue of stellar spectra relevant for astrophysical applications
3.1 Independent Gaussian distributions dataset
@@
The first dataset is structured with 2 target classes and input vectors with 20 independent features. Each data component is sampled from a Gaussian distribution, see Fig. 2. While some components are drawn from distributions which are set to be identical across classes, others display only partial overlap. Components with identical distributions offer no discriminatory information. As such, they prove ineffective for classification purposes. Conversely, the other components are pivotal, the smaller the overlap the larger their relevance in promoting the sought discrimination.
Figure 2: Input vector components distributions. Distribution of the values of the 20 components of the input vectors. The distributions for the data belonging to the first (resp. second) class are represented in red (resp. blue). For some components, the two classes are indistinguishable, while in some other cases the two distributions are only partially overlapping.
@@
A spectral feedforward neural network, parametrized as stipulated by Eq. (1), has been trained with the aim of assigning input from the above dataset to the correct class of origin. The employed network consists of three hidden layers with, respectively, 100, 100, and 50 neurons. In addition, the network accommodates for an input layer of size 20 and an output layer of size 2. At the end of the training, we analyzed the 20 eigenvalues, associated with the input layer as in the spirit of the above discussion (with an abuse of language we will from here on call eigenvalues the rescaled version of this latter which incorporate the norm of the eigenvectors, as stipulated in the Methods section). As shown in the left panel of Fig. 3, eigenvalues entangled with irrelevant features are all zero, while eigenvalues which can be traced back to relevant input components are non-zero. Hence, the eigenvalues associated to the input layer can efficiently tag features identified as relevant by the network itself to solve the assigned task. The magnitude of non-zero eigenvalues, i.e., those associated with relevant components, is found to correlate with a direct measure of the relative distance between the distributions from which individual components are drawn (see the right panel of Fig. 3). Since the distributions differ only in their mean value, we quantify the above distance as the absolute value of the difference between the two means.
@@
Figure 3: The left panel displays the 20 post-training eigenvalues associated to the input layer, for the simple Gaussian dataset. The red dots and the red bars correspond to the average values and the standard deviations of the results obtained by repeating the experiments for 10 independent realizations. In the right panel, the same post-training eigenvalues are reported against a relevance parameter that gauges the distance between the distributions (respectively blue and red, in Fig. 2) from which the components are eventually drawn. The distance is defined as the absolute value of the difference between the two means.
3.2 Correlated Gaussian distributions dataset
@@
Unlike the dataset just analyzed, real datasets often exhibit correlations between input vector components, and crucial task-solving information are frequently concealed within those correlations. With this motivation in mind, we tested the proposed spectral feature detection method on a different dataset that contains pairwise correlations between input vector components. The dataset is still organized into two distinct classes that are to be eventually identified. Input vectors are now made of 10 components. The first 8 components are defined as follows:
$\begin{split}&x_{2n}\sim\mathcal{N}(0,1)\\ &x_{2n+1}=x_{2n}\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\text{if}\;\;c(\vec{x})=0\\ &x_{2n+1}=\begin{cases}x_{2n}\;\text{with prob.}\;1-p_{n}\\ -x_{2n}\;\text{with prob.}\;p_{n}\end{cases}\;\;\text{if}\;\;c(\vec{x})=1\end{split}$ (2)
@@
where $c(\vec{x})$ is the target class of the input and the index $n\in\{0,1,2,3\}$ denotes the four pairs. The last two components are independently drawn from two identical normal distributions $\mathcal{N}(1,0)$ regardless of the class of belonging:
@@
As a first noticeable observation, we remark that it is not possible to separate the two datasets by solely relying on independent observations of the respective vector components. Indeed, crucial information are reflected in how pairs of components are mutually correlated (see left panels of Fig. 4).
@@
To better understand the relationship between the different components, let us focus on the first two entries of the input vectors, $x_{0}$ and $x_{1}$. For data belonging to class 00 (as we label the first of the two) $x_{0}=x_{1}$, that is, the first two components are perfectly correlated. Conversely, the two components of data belonging to the other class (denoted by $1$) are equal with probability $1-p_{0}$ and have opposite sign with probability $p_{0}$. If $p_{0}=0$, the first two components are equal for both classes. No information about the class to which the supplied data belong can be thus extracted by comparing the two above components. The second component is indeed a mere copy of the first and therefore irrelevant for classification purposes. On the other hand, if $p_{0}=1$, the first and second components from class $1$ data are in perfect anti-correlation, while data from class 00 are still matched and thus fully correlated. Hence, in the latter case, it is possible to completely separate the two classes by solely comparing $x_{0}$ and $x_{1}$. The Parameter $p_{n}$ quantifies therefore the degree of sensible information which can be pulled out from a direct comparison of the two components in the $n$-th pair. We want to stress that, regardless of the value of $p_{n}$, neither component individually contains information on the specific class to which the vector belongs. Therefore, useful information can be obtained to solve the classification task only by analyzing the two components simultaneously. Following this observations, we generated the dataset by using four different values of $p_{n}\in[0,1]$ for the four pairs. Finally, the last two components are not correlated with the others and are identically distributed for both classes, hence void of information. We therefore assign a parameter $p_{4}=0$ to these two components to explicitly state their manifest irrelevance.
@@
Figure 4: Left column: the distributions of the values of the fourth and fifth components of the input vectors in the correlated Gaussian dataset are shown in the first two panels. In the bottom-left panel, the fifth component is plotted against the corresponding fourth component for some input vectors of the same dataset. Right column: the upper-right panel reports the ten post-training eigenvalues (refereed to the input layer) for the correlated Gaussian dataset. Red dots and red bars quantifies the recorded average values and the associated standard deviations, as obtained by averaging over 10 independent realizations. In the bottom-right panel the same post-training eigenvalues are plotted against the parameter $p_{n}$, as defined in the main text.
@@
The architecture of the employed neural network, parameterized via its spectral modality, is identical to that put forward for the analysis of the uncorrelated dataset (except for the first layer that in this case receives just 10 entries). In upper right panel of Fig. 4, the post-training eigenvalues referred to the input layer are shown. Firstly, we notice that pairs are correctly spotted out, as they are assigned the very same eigenvalue entry. Moreover, eigenvalues associated with irrelevant components are tuned zero for this dataset as well, an observation that testifies again on the ability of the spectral network to discriminate between irrelevant and relevant features. Lastly, the magnitude of the recorded eigenvalues scales proportionally to the control parameter $p_{n}$, which quantifies the relevance of the selected pair for classification purposes (bottom right panel, Fig. 4). In conclusion, and also for this second, more complex dataset, the optimized eigenvalues referred to the first layer are a good proxy of the importance of the associated components.
3.3 The MNIST dataset